OpenStudy (anonymous):

evaluate the expression log[2] 16-log[5] 125

OpenStudy (anonymous):

log 2 16 = 4 log 5 125 = 3 4 + 3 = 7 .. the answer

OpenStudy (anonymous):

4 - 3 = 1

OpenStudy (anonymous):

\[\log_{2 } 16 - \log_{5} 125\]\[\log_{2} 4^2 - \log_{5} 5^3\]\[4\log_{2} 2 - 3\log_{5} 5\]4-3 = 1